OpenStudy (anonymous):

2x-11 square root x +12=0 solve. please and Thank you

OpenStudy (compassionate):

\[2x-11\sqrt{x + 12} = 0\]

OpenStudy (compassionate):

^ Is that correct?

OpenStudy (anonymous):

no it's \[2x-11+\sqrt{x}+12=0\]

OpenStudy (anonymous):

See, you can subtract 11 from 12: you will get: \[2x + \sqrt{x} + 1 = 0 \implies (2x + 1) = -\sqrt{x}\] Now squaring both the sides you will get: \[(2x + 1)^2 = (-\sqrt{x})^2 \implies 4x^2 + 1 + 4x = x\] \[4x^2 + 3x + 1 = 0\] Use quadratic Formula to find the values of x..

OpenStudy (anonymous):

Thank you